OpenStudy (anonymous):

Numbers 1 through 10 are written on cards and placed in a bag. Find the probability of choosing a number greater than 5 or choosing an odd number

OpenStudy (anonymous):

The probability of a number greater then five is 5/10 The probability of a number that is odd is also 5/10 the answer is \[\frac{ 5 }{ 10 }\]

OpenStudy (anonymous):

@catkittycool321 the only options they give me are 1/5 , 2/5, 3/5, or 4/5

OpenStudy (johnweldon1993):

So there are a total of 10 cards.....And you want either a number > 5 or an odd number there are 5 numbers > 5 (6,7,8,9,10) and there are 5 numbers that are odd (1,3,5,7,9) you want the probability of choosing either or....so 5 are > 5 3 more are odd so You have 8 possibilities that you want \[\frac{ 8 }{ 10 }\] simplify this down to get \[\frac{ 4 }{ 5 }\]

OpenStudy (mathstudent55):

Total number of cards? 10 Cards with number greater than 5? 6, 7, 8, 9, 10, or 5 cards with desired outcomes. Cards with odd numbers? 1, 3, 5, 7, 9, of which 1, 3, and 5 are not included in numbers greater than 5. So this adds 3 more desired oucomes. Total number of desired outcomes: 8 Probability = desired outcomes/total outcomes = 8/10 = 4/5

OpenStudy (anonymous):

thank you guys!

OpenStudy (johnweldon1993):

No problem!

OpenStudy (mathstudent55):

wlcm